MICROWAVE SPECTRA AND NUCLEAR QUADRUPOLE COUPLING IN F_{3}B-N(CH_{3})_{3} AND (CH_{3})_{3}B-N(CH_{3})_{3}

D. L. FIACCO, \underline{S. W. HUNT}, and K. R. LEOPOLD, Department of Chemistry, University of Minnesota, Minneapolis, MN 55455.

The donor-acceptor complexes F_{3}B-N(CH_{3})_{3} and (CH_{3})_{3}B-N(CH_{3})_{3} have been studied by Fourier transform microwave spectroscopy. The rotational constants for both species are consistent with the literature values. For F_{3}B-N(CH_{3})_{3}, the J = 0 - 1 and 1 - 2 transitions have been fully analyzed for the ^{1}^{0}B, ^{1}^{1}B, ^{1}^{4}N and ^{1}^{5}N species. A single state is observed, with no evidence of internal rotation. For (CH_{3})_{3}B-N(CH_{3})_{3}, the 0-1 and 1-2 transitions have been recorded and analyzed. In this case, two states are observed, with an intensity ratio of about 10:1. For both systems, a Townes and Dailey analysis of the ^{1}^{4}N nuclear quadrupole coupling constants has been used to provide information about the electronic structure of the complexes. In the case of F_{3}B-N(CH_{3})_{3}, about 0.6 electrons are transferred from the trimethylamine to the BF_{3} upon formation of the donor-acceptor bond, while for (CH_{3})_{3}B-N(CH_{3})_{3} a value of 0.42 electrons is obtained. These results will be compared with similar measurements for related adducts of ammonia and trimethylamine.